OpenStudy (summersnow8):

MOPS is often used as a buffer system in biochemistry. What is the pH of a buffer solution prepared by combining 0.200 moles of MOPS base and 0.120 moles of HCl in 2.00 L of water? (pKa = 7.20 for MOPS-H+) The answer is 7.02 MOPS-Na + HCL → MOPS-H + NaCl

OpenStudy (photon336):

Hey how has your class been so far? I've been away from OS for sometime.

OpenStudy (photon336):

pH = pkA+Log[[A]/HA]

OpenStudy (jfraser):

@Photon336 has given you the Henderson-Hasselbalch equation, which combines the pieces to make a useful buffer. A good buffer must have a weak acid (MOPS-H), and its conjugate (MOPS-), and both in approximately equal amounts if it's going to be effective. The ratio of those two concentrations will change the pH of the buffer + or - relative to its pKa, depending on which component is there in greater concentrations.

OpenStudy (summersnow8):

@JFraser so when I plug in the numbers to the equation: pH = pKa + log (base/acid) pH = 7.20 + log (0.200/0.120) pH = 7.42, which is too high

OpenStudy (jfraser):

you've plugged in the \(moles\) of each, you need to find the \(concentration\) of each

OpenStudy (summersnow8):

well the concentration of HCl can be found by multiplying moles by L: 2.00 L * 0.120 moles = 0.24 M but I don't know about the buffer

OpenStudy (jfraser):

but you don't actually have the moles of MOPS-H, you're given the moles of HCl. You need to do the stoichiometry first to find out how much MOPS-H you form

OpenStudy (summersnow8):

Like this: MOPS-Na + HCL → MOPS-H + NaCl 0.200 moles 0.120 moles x

OpenStudy (jfraser):

basically, yes

OpenStudy (jfraser):

how many moles of MOPS- are \(left over\), and how many moles of MOPS-H are \(formed\)?

OpenStudy (summersnow8):

I am not sure how you figure that out

OpenStudy (jfraser):

this is a limiting reactant problem. If you start with 0.200mol of MOPS-, and 0.120mol of HCl, \(all\) of the HCl will be used, but only \(some\) of the MOPS- will be used, won't it?

OpenStudy (summersnow8):

moles of MOPS-H = 0.32 moles of MOPS left over is 0.32 - 0.200 = 0.12

OpenStudy (jfraser):

no, the moles of MOPS-H formed is .0.120

OpenStudy (summersnow8):

How did you determine that?

OpenStudy (jfraser):

because the products are formed, as the reactants are used up

OpenStudy (summersnow8):

MOPS-Na + HCL → MOPS-H + NaCl 0.200 moles 0.120 moles x - 0.120 moles -0.120 moles + 0.120 moles ---------------------------------------------- 0.08 moles 0 0.120 moles

OpenStudy (summersnow8):

is that what you mean?

OpenStudy (summersnow8):

MOPS left over is 0.08 moles MOPS-H is 0.120 moles

OpenStudy (jfraser):

exactly, yes

OpenStudy (jfraser):

now find the concentrations of MOPS- (that's A-) and MOPS-H (that's HA) and solve

OpenStudy (summersnow8):

well to find concentration I know you want to multiply L by moles, but only 2.00 L of water was added

OpenStudy (summersnow8):

I am not sure where to go to from here.... I thought I would multiply both by 2.00 L, but that doesn't work

OpenStudy (jfraser):

concentration is moles \(per\) liter, so you \(divide\) by the total volume of 2.00L

OpenStudy (summersnow8):

so, 0.08 moles / 2.00 L = 0.04 M 0.120 moles / 2.00 L = 0.06 M pH = 7.20 + log (0.04 / 0.06) pH = 7.02

OpenStudy (jfraser):

it fits, GJ